Calculate the slope of a line passing through (x,5) and (7,2)
Answer by oberobic(2304)

(x,y) = (x,5)
.
(x1,y1) = (7,2)
.
m = (y - y1) / (x - x1)
.
m = (5-7) / (x-2)
.
m = -2 / (x-2)
.
Done.